# NeuralSurv: Deep Survival Analysis with Bayesian Uncertainty Quantification

### v1

@@ G.4 Optimal Update for $\boldsymbol{\theta}$
Using standard mean-field variational inference techniques (see, e.g., Chapter 10.1 of [3]), the optimal variational factor for the parameters $\boldsymbol{\theta}$ is obtained by computing the expectation of the joint log-density with respect to the other variational factors. In particular, we have
@@
where $\lambda^{\mathbb{Q},(k)}_{i}(t,\omega)$ is shown in Equation (40). Next, we recall the expression for $g^{\text{lin}}(\cdot;\boldsymbol{\theta})$ from (17) and we notice that
$g^{\text{lin}}(\cdot;\boldsymbol{\theta})$ $=\boldsymbol{\theta}^{\top}\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(\cdot)+\text{const.}$.
$g^{\text{lin}}(\cdot;\boldsymbol{\theta})^{2}$ $=\boldsymbol{\theta}^{\top}\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(\cdot)\left(2g(\cdot;\boldsymbol{\theta}_{\text{MAP}})-2\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(\cdot)^{\top}\boldsymbol{\theta}_{\text{MAP}}\right)+\boldsymbol{\theta}^{\top}\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(\cdot)\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(\cdot)^{\top}\boldsymbol{\theta}+\text{const.},$. ,
@@
$\quad-\left(\mathcal{I}^{(k)}_{1,i}+2\left(\mathcal{I}^{(k)}_{2,i}-\mathcal{I}_{3,i}^{(k)}\boldsymbol{\theta}_{\text{MAP}}\right)\right)\Bigg{)}$ (43)
$\mathbf{B}^{(k)}$ $=\sum_{i=1}^{N}\frac{1}{2}\left(\delta_{i}\>\mathbb{E}_{\omega_{i}\sim q^{(k)}_{\omega_{i}}}[\omega_{i}]\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(y_{i},\mathbf{x}_{i})\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(y_{i},\mathbf{x}_{i})^{\top}+\mathcal{I}^{(k)}_{3,i}\right)+\frac{1}{2}\mathbf{I}_{m}$ (44)
@@
$\tilde{\boldsymbol{\mu}}^{(k)}=\frac{1}{2}\left(\mathbf{B}^{(k)}\right)^{-1}\mathbf{A}^{(k)},\quad\tilde{\boldsymbol{\Sigma}}=\frac{1}{2}\left(\mathbf{B}^{(k)}\right)^{-1}.$ (45)
@@
$\mathbb{E}_{\boldsymbol{\theta}\sim q^{(k)}_{\boldsymbol{\theta}}}[g^{\text{lin}}(t,\mathbf{x}_{i};\boldsymbol{\theta})]$ $=g(t,\mathbf{x}_{i};\boldsymbol{\theta}_{\text{MAP}})+\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(t,\mathbf{x}_{i})^{\top}\left(\tilde{\boldsymbol{\mu}}^{(k)}-\boldsymbol{\theta}_{\text{MAP}}\right),$ (46)
@@ Appendix H Coordinate Ascent Variational Inference Algorithm
1:  Compute: Compute $\tilde{\beta}$ following (41).
@@
3:  Compute: $\mathbb{E}_{\phi\sim q^{(0)}_{\phi}}[\log\phi]$ given $\left(\tilde{\alpha}^{(0)},\tilde{\beta}\right)$ following (LABEL:eq:optimal_variational_update_phi_gamma_expectation).
4:  Compute: $\left\{(\tilde{m}_{i}(\cdot),\tilde{s}_{i}(\cdot))^{(0)}\right\}_{i=1}^{N}$ given $\left(\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\Sigma}}\right)^{(0)}$ following (46).
@@
9:         Update: $\left\{\tilde{c}_{i}^{(k)}\right\}_{i=1}^{N}$ given $\left\{\tilde{s}_{i}(\cdot)^{(k-1)}\right\}_{i=1}^{N}$ following (35).
10:         Compute: $\left\{\mathbb{E}_{\omega_{i}\sim q^{(k)}_{\omega_{i}}}[\omega_{i}]\right\}_{i=1}^{N}$ given $\left\{\tilde{c}_{i}^{(k)}\right\}_{i=1}^{N}$ following (36).
@@
12:         Update: $\left\{\lambda^{\mathbb{Q},(k)}_{i}(\cdot)\right\}_{i=1}^{N}$ given ($\left(\{(\tilde{m}_{i}(\cdot),\tilde{s}_{i}(\cdot))^{(k-1)}\}_{i=1}^{N},\mathbb{E}_{\phi\sim q^{(k-1)}_{\phi}}[\log\phi]\right)$ following (40).
@@
14:         Update: $\tilde{\alpha}^{(k)}$ given $\left\{\lambda^{\mathbb{Q},(k)}_{i}(\cdot)\right\}_{i=1}^{N}$ following (41).
15:         Compute: $\mathbb{E}_{\phi\sim q^{(k)}_{\phi}}[\log\phi]$ given $(\tilde{\alpha}^{(k)},\tilde{\beta})$ following (LABEL:eq:optimal_variational_update_phi_gamma_expectation).
@@
17:         Update: $\left(\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\Sigma}}\right)^{(k)}$ given $\left\{(\mathbb{E}_{\omega_{i}\sim q^{(k)}_{\omega_{i}}}[\omega_{i}],\lambda^{\mathbb{Q},(k)}_{i}(\cdot)\right\}_{i=1}^{N}$ following (45).
18:         Compute: $\left\{(\tilde{m}_{i}(\cdot),\tilde{s}_{i}(\cdot))^{(k)}\right\}_{i=1}^{N}$ given $\left(\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\Sigma}}\right)^{(k)}$ following (46).
@@
20:  Return: Optimized variational distributions $q^{(k^{\star})}_{\boldsymbol{\theta}}(\boldsymbol{\theta})=\mathcal{N}\left(\tilde{\boldsymbol{\mu}}^{(k^{\star})},\tilde{\boldsymbol{\Sigma}}^{(k^{\star})}\right)$ and $q^{k^{\star}}_{\phi}(\phi)=\text{Gamma}\left(\tilde{\alpha}^{(k^{\star})},\tilde{\beta}\right)$, where $k^{\star}$ is the final iteration after convergence.
@@ Appendix I Computational Speed-Ups
for $i=1,\ldots,N$. With this notation (and dropping the CAVI-iteration index for clarity), the matrix $\mathbf{B}\in\mathbb{R}^{m\times m}$ defined in (G.4) becomes
@@
and in general admits no closed-form solution. We therefore approximate it by any standard quadrature rule (e.g. trapezoid, Simpson’s, or Gauss–Legendre). In what follows, we will illustrate the argument with the trapezoid rule, though the same steps apply to any other quadrature method.
@@
where $t_{1}:=0$ and $t_{K}:=\max\{y_{i}\}_{i=1}^{N}$.We associate a set of quadrature weights $\{v_{ik}\}_{k=1}^{K}$ to the time grid points, tailored for each observation $i$. These weights correspond to the trapezoidal rule for numerical integration on the interval $[0,y_{i}]$, and are defined as:
@@
$\mathbf{Q}_{i}$ $:=\begin{bmatrix}\mathbf{J}_{\boldsymbol{\theta}_{\mathrm{MAP}}}(t_{1},\mathbf{x}_{i})&\mathbf{J}_{\boldsymbol{\theta}_{\mathrm{MAP}}}(t_{2},\mathbf{x}_{i})&\cdots&\mathbf{J}_{\boldsymbol{\theta}_{\mathrm{MAP}}}(t_{K},\mathbf{x}_{i})\end{bmatrix}\;\in\;\mathbb{R}^{m\times K}.$
@@
Applying the Woodbury identity (see [16, Appendix B.10]) then reduces the inversion of $\mathbf{B}$ to that of an $R\times R$ matrix:
@@
instead of $\mathcal{O}(m^{3})$ for a full $m\times m$ inversion. Whenever $R\ll m$, this yields a dramatic speed-up. By replacing the direct $\mathcal{O}(R^{3})$ factorization with a Conjugate-Gradient (CG) solver — as is commonly done in Gaussian-process toolkits such as GPyTorch [11] — we reduce the cost to $\mathcal{O}(R^{2})$.
@@
The first successful trials of adjuvant chemotherapy for colon cancer dataset was obtained from the survival package [39]. The dataset contains records of 1,822 observations with 15 covariates among which 49.23% are censored. All rows with missing values were excluded from the dataset.
@@
The National Wilm’s Tumor Study (NWTCO) was obtained from the pycox package [23]. The dataset contains records of 4,028 observations with 7 covariates among which 14.18% are censored.
@@
The Rotterdam and German Breast Cancer Study Group (GBSG) was obtained from the pycox package [23]. The dataset contains records of 2,232 observations with 7 covariates among which 43.23% are censored.
@@
The Molecular Taxonomy of Breast Cancer International Consortium (METABRIC) dataset was obtained from the pycox package [23]. The dataset contains records of 1,904 observations with 9 covariates among which 42.07% are censored.
@@
The Worcester Heart Attack Study (WHAS) dataset was obtained from the sksurv package [35]. The dataset contains records of 500 observations with 14 covariates among which 43.00% are censored.
@@
The Study to Understand Prognoses and Preferences for Outcomes and Risks of Treatment (SUPPORT) dataset was obtained from the pycox package [23]. The dataset contains records of 8,873 observations with 14 covariates among which 31.97% are censored.
@@
The Veterans administration Lung Cancer trial (VLC) dataset was obtained from the sksurv package [35]. The dataset contains records of 137 observations with 8 covariates among which 6.57% are censored.
@@
The Sac 3 dataset from the simulation study in [24, Appendix A.1] was obtained from the pycox package [23]. The dataset contains records of 100,000 observations with 45 covariates among which 37.20% are censored.
@@
The Multi-Task Logistic Regression [43] was implemented using the MTLR class from the pycox package [23].
@@
The DeepHit method [28] was implemented using the DeepHitSingle class from the pycox package [23]. The hyperparameters $\alpha$ and $\sigma$ were set to 0.2 and 0.1, respectively. Those are the default values.
@@
The DeepSurv model [20] was implemented using the CoxPH class from the pycox package [23].
@@
The Logistic Hazard method [43] was implemented using the LogisticHazard class from the pycox package [23].
@@
The CoxTime method [25] was implemented using the CoxTime class from the pycox package [23].
@@
The CoxCC method [25] was implemented using the CoxCC class from the pycox package [23].
@@
The PMF method [24] was implemented using the PMF class from the pycox package [23].
@@
The PCHazard method [24] was implemented using the PCHazard class from the pycox package [23].
@@
The BCESurv method [24] was implemented using the BCESurv class from the pycox package [23].
@@
The DySurv method [32] was implemented using the official code provided by the authors, available at https://github.com/munibmesinovic/DySurv/blob/main/Models/Results/Static_Benchmarks_GBSG_Example.ipynb (Accessed on May 13 2025).
@@
The Cox Proportional Hazards model [7] was implemented using the CoxPHFitter class from the lifelines package [8]. The Breslow method was used to compute the survival function.
@@
The Weibull Accelerated Failure Time model [6] was implemented using the WeibullAFTFitter class from the lifelines package [8].
@@
The Random Survival Forest [19] was implemented using the RandomSurvivalForest class from the sksurv package [35]. The number of trees in the forest is set to 1,000. The minimum number of samples required to split an internal node is 10, and the minimum number of samples required to be at a leaf node is 15. Those were the same hyperparameters as used in [32].
@@
The Survival Support Vector Machine [36] was implemented using the FastSurvivalSVM class from the sksurv package [35]. The optimal regularization hyperparameter $\alpha$ was selected via grid search by evaluating model performance on the training set using the C-index. This method does not allow for estimation of the survival function. Predicted ranks were used as risk scores for computing the C-index.
@@
Let $\hat{q}_{i}(t)$ be the predicted risk score of observation with covariates $\mathbf{x}_{i}$ at time $t$. The C-index estimate [15] is given by
[…]
Let $\hat{S}_{i}(t)$ be the predicted survival function of observation with covariates $\mathbf{x}_{i}$ at time $t$. When the predicted risk score is taken to be the negative of the survival function, i.e., $\hat{q}_{i}(t)=-\hat{S}_{i}(t)$, the C-index is referred to as the Antolini’s C-index [1] and is found with
[…]
The C-index is obtained using the ConcordanceIndex class from the TorchSurv package [33].
IPCW Integrated Brier Score
Let $\hat{S}_{i}(t)$ be the predicted survival function of observation with covariates $\mathbf{x}_{i}$ at time $t$. Let the inverse probability censoring weight (IPCW) at time $t$ be defined as the inverse of the probability of being uncensored, $\xi(t)=1/\hat{C}(t)$, where $\hat{C}(t)$ denotes the Kaplan–Meier estimate of the censoring survival function. Under right censorship, the IPCW Brier score (BS) [14] at time $t$ is given by
IPCW $\text{IPCW BS}(t)=\frac{1}{N}\sum_{i=1}^{N}\xi(y_{i})\mathbbm{1}_{\{y_{i}\leq t,\delta_{i}=1\}}(0-\hat{S}_{i}(t))^{2}+\xi(t)\mathbbm{1}_{\{y_{i}>t\}}(1-\hat{S}_{i}(t))^{2}.$ (47)
The IBS is the integral of the Brier Score in (47). The IPCW weights and the IPCW IBS are computed using the get_ipcw function and the BrierScore class from the TorchSurv package [33].
@@
The parameters are initialized so that they match their prior expected values. Specifically, we set $\boldsymbol{\theta}^{(0)}=\mathbf{0}$ and $\phi^{(0)}=\alpha_{0}/\beta_{0}$ The maximization step of the EM algorithm is performed using the L-BFGS-B algorithm. The EM algorithm is considered to have converged when the relative change in the Q-function between consecutive iterations falls below a tolerance threshold of $10^{-6}$ for two successive iterations.

### v2

@@ G.4 Optimal Update for $\boldsymbol{\theta}$
Using standard mean-field variational inference techniques (see, e.g., Chapter 10.1 of [bishop_pattern_recognition]), the optimal variational factor for the parameters $\boldsymbol{\theta}$ is obtained by computing the expectation of the joint log-density with respect to the other variational factors. In particular, we have
@@
where $\lambda^{\mathbb{Q},(k)}_{i}(t,\omega)$ is shown in Equation (A22). Next, we recall the expression for $g^{\text{lin}}(\cdot;\boldsymbol{\theta})$ from (18) and we notice that
$g^{\text{lin}}(\cdot;\boldsymbol{\theta})$ $=\boldsymbol{\theta}^{\top}\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(\cdot)+\text{const.}$
$g^{\text{lin}}(\cdot;\boldsymbol{\theta})^{2}$ $=\boldsymbol{\theta}^{\top}\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(\cdot)\left(2g(\cdot;\boldsymbol{\theta}_{\text{MAP}})-2\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(\cdot)^{\top}\boldsymbol{\theta}_{\text{MAP}}\right)+\boldsymbol{\theta}^{\top}\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(\cdot)\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(\cdot)^{\top}\boldsymbol{\theta}+\text{const.},$
@@
$\quad-\left(\mathcal{I}^{(k)}_{1,i}+2\left(\mathcal{I}^{(k)}_{2,i}-\mathcal{I}_{3,i}^{(k)}\boldsymbol{\theta}_{\text{MAP}}\right)\right)\Bigg)$ (A25)
$\mathbf{B}^{(k)}$ $=\sum_{i=1}^{N}\frac{1}{2}\left(\delta_{i}\>\mathbb{E}_{\omega_{i}\sim q^{(k)}_{\omega_{i}}}[\omega_{i}]\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(y_{i},\mathbf{x}_{i})\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(y_{i},\mathbf{x}_{i})^{\top}+\mathcal{I}^{(k)}_{3,i}\right)+\frac{1}{2}\mathbf{I}_{m}$ (A26)
@@
$\tilde{\boldsymbol{\mu}}^{(k)}=\frac{1}{2}\left(\mathbf{B}^{(k)}\right)^{-1}\mathbf{A}^{(k)},\quad\tilde{\boldsymbol{\Sigma}}=\frac{1}{2}\left(\mathbf{B}^{(k)}\right)^{-1}.$ (A27)
@@
$\mathbb{E}_{\boldsymbol{\theta}\sim q^{(k)}_{\boldsymbol{\theta}}}[g^{\text{lin}}(t,\mathbf{x}_{i};\boldsymbol{\theta})]$ $=g(t,\mathbf{x}_{i};\boldsymbol{\theta}_{\text{MAP}})+\mathbf{J}_{\boldsymbol{\theta}_{\text{MAP}}}(t,\mathbf{x}_{i})^{\top}\left(\tilde{\boldsymbol{\mu}}^{(k)}-\boldsymbol{\theta}_{\text{MAP}}\right),$ (A28)
@@ Appendix H Coordinate Ascent Variational Inference Algorithm
1:Compute: Compute $\tilde{\beta}$ following (A23).
@@
[…]
4:Compute: $\left\{(\tilde{m}_{i}(\cdot),\tilde{s}_{i}(\cdot))^{(0)}\right\}_{i=1}^{N}$ given $\left(\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\Sigma}}\right)^{(0)}$ following (A28).
@@
9:    Update: $\left\{\tilde{c}_{i}^{(k)}\right\}_{i=1}^{N}$ given $\left\{\tilde{s}_{i}(\cdot)^{(k-1)}\right\}_{i=1}^{N}$ following (A17).
10:    Compute: $\left\{\mathbb{E}_{\omega_{i}\sim q^{(k)}_{\omega_{i}}}[\omega_{i}]\right\}_{i=1}^{N}$ given $\left\{\tilde{c}_{i}^{(k)}\right\}_{i=1}^{N}$ following (A18).
@@
12:    Update: $\left\{\lambda^{\mathbb{Q},(k)}_{i}(\cdot)\right\}_{i=1}^{N}$ given $\left(\{(\tilde{m}_{i}(\cdot),\tilde{s}_{i}(\cdot))^{(k-1)}\}_{i=1}^{N},\mathbb{E}_{\phi\sim q^{(k-1)}_{\phi}}[\log\phi]\right)$ following (A22).
@@
14:    Update: $\tilde{\alpha}^{(k)}$ given $\left\{\lambda^{\mathbb{Q},(k)}_{i}(\cdot)\right\}_{i=1}^{N}$ following (A23).
[…]
@@
17:    Update: $\left(\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\Sigma}}\right)^{(k)}$ given $\left\{(\mathbb{E}_{\omega_{i}\sim q^{(k)}_{\omega_{i}}}[\omega_{i}],\lambda^{\mathbb{Q},(k)}_{i}(\cdot)\right\}_{i=1}^{N}$ following (A27).
18:    Compute: $\left\{(\tilde{m}_{i}(\cdot),\tilde{s}_{i}(\cdot))^{(k)}\right\}_{i=1}^{N}$ given $\left(\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\Sigma}}\right)^{(k)}$ following (A28).
@@
[…]
@@ Appendix I Computational Speed-Ups
for $i=1,\ldots,N$. With this notation (and dropping the CAVI‐iteration index for clarity), the matrix $\mathbf{B}\in\mathbb{R}^{m\times m}$ defined in (G.4) becomes
@@
and in general admits no closed‐form solution. We therefore approximate it by any standard quadrature rule (e.g. trapezoid, Simpson’s, or Gauss–Legendre). In what follows, we will illustrate the argument with the trapezoid rule, though the same steps apply to any other quadrature method.
@@
where $t_{1}:=0$ and $t_{K}:=\max\{y_{i}\}_{i=1}^{N}$.We associate a set of quadrature weights $\{v_{ik}\}_{k=1}^{K}$ to the time grid points, tailored for each observation $i$. These weights correspond to the trapezoidal rule for numerical integration on the interval $0,y_{i}]$, and are defined as:
@@
$\mathbf{Q}_{i}$ $=\begin{bmatrix}\mathbf{J}_{\boldsymbol{\theta}_{\mathrm{MAP}}}(t_{1},\mathbf{x}_{i})&\mathbf{J}_{\boldsymbol{\theta}_{\mathrm{MAP}}}(t_{2},\mathbf{x}_{i})&\cdots&\mathbf{J}_{\boldsymbol{\theta}_{\mathrm{MAP}}}(t_{K},\mathbf{x}_{i})\end{bmatrix}\;\in\;\mathbb{R}^{m\times K}.$
@@
Applying the Woodbury identity (see [Higham_matrices, Appendix B.10]) then reduces the inversion of $\mathbf{B}$ to that of an $R\times R$ matrix:
@@
instead of $\mathcal{O}(m^{3})$ for a full $m\times m$ inversion. Whenever $R\ll m$, this yields a dramatic speed-up. By replacing the direct $\mathcal{O}(R^{3})$ factorization with a Conjugate-Gradient (CG) solver — as is commonly done in Gaussian-process toolkits such as GPyTorch [gpytorch] — we reduce the cost to $\mathcal{O}(R^{2})$.
@@
The first successful trials of adjuvant chemotherapy for colon cancer dataset was obtained from the survival package [survivalpackage]. The dataset contains records of 1,822 observations with 15 covariates among which 49.23% are censored. All rows with missing values were excluded from the dataset.
@@
The National Wilm’s Tumor Study (NWTCO) was obtained from the pycox package [pycox]. The dataset contains records of 4,028 observations with 7 covariates among which 14.18% are censored.
@@
The Rotterdam and German Breast Cancer Study Group (GBSG) was obtained from the pycox package [pycox]. The dataset contains records of 2,232 observations with 7 covariates among which 43.23% are censored.
@@
The Molecular Taxonomy of Breast Cancer International Consortium (METABRIC) dataset was obtained from the pycox package [pycox]. The dataset contains records of 1,904 observations with 9 covariates among which 42.07% are censored.
@@
The Worcester Heart Attack Study (WHAS) dataset was obtained from the sksurv package [sksurv]. The dataset contains records of 500 observations with 14 covariates among which 43.00% are censored.
@@
The Study to Understand Prognoses and Preferences for Outcomes and Risks of Treatment (SUPPORT) dataset was obtained from the pycox package [pycox]. The dataset contains records of 8,873 observations with 14 covariates among which 31.97% are censored.
@@
The Veterans administration Lung Cancer trial (VLC) dataset was obtained from the sksurv package [sksurv]. The dataset contains records of 137 observations with 8 covariates among which 6.57% are censored.
@@
The Sac 3 dataset from the simulation study in [Kvamme2021, Appendix A.1] was obtained from the pycox package [pycox]. The dataset contains records of 100,000 observations with 45 covariates among which 37.20% are censored.
@@
The Multi-Task Logistic Regression [Yu2011] was implemented using the MTLR class from the pycox package [pycox].
@@
The DeepHit method [Lee2018] was implemented using the DeepHitSingle class from the pycox package [pycox]. The hyperparameters $\alpha$ and $\sigma$ were set to 0.2 and 0.1, respectively. Those are the default values.
@@
The DeepSurv model [Katzman2018] was implemented using the CoxPH class from the pycox package [pycox].
@@
The Logistic Hazard method [Yu2011] was implemented using the LogisticHazard class from the pycox package [pycox].
@@
The CoxTime method [Kvamme2019] was implemented using the CoxTime class from the pycox package [pycox].
@@
The CoxCC method [Kvamme2019] was implemented using the CoxCC class from the pycox package [pycox].
@@
The PMF method [Kvamme2021] was implemented using the PMF class from the pycox package [pycox].
@@
The PCHazard method [Kvamme2021] was implemented using the PCHazard class from the pycox package [pycox].
@@
The BCESurv method [Kvamme2021] was implemented using the BCESurv class from the pycox package [pycox].
@@
The DySurv method [Mesinovic2024] was implemented using the official code provided by the authors, available at https://github.com/munibmesinovic/DySurv/blob/main/Models/Results/Static_Benchmarks_GBSG_Example.ipynb (Accessed on May 13 2025).
Sumo-Net.
The Sumo-Net method [rindt2022] was implemented using the official code provided by the authors, available at https://github.com/MrHuff/Sumo-Net (Accessed on July 25 2025).
DQS.
The DQS method [Yanagisaw2023] was implemented using the official code provided by the authors, available at https://github.com/IBM/dqs (Accessed on July 25 2025).
@@
The Cox Proportional Hazards model [Cox1972] was implemented using the CoxPHFitter class from the lifelines package [davidson2019lifelines]. The Breslow method was used to compute the survival function.
@@
The Weibull Accelerated Failure Time model [Carroll2003] was implemented using the WeibullAFTFitter class from the lifelines package [davidson2019lifelines].
@@
The Random Survival Forest [Ishwaran2008] was implemented using the RandomSurvivalForest class from the sksurv package [sksurv]. The number of trees in the forest is set to 1,000. The minimum number of samples required to split an internal node is 10, and the minimum number of samples required to be at a leaf node is 15. Those were the same hyperparameters as used in [Mesinovic2024].
@@
The Survival Support Vector Machine [Polsterl2015] was implemented using the FastSurvivalSVM class from the sksurv package [sksurv]. The optimal regularization hyperparameter $\alpha$ was selected via grid search by evaluating model performance on the training set using the C-index. This method does not allow for estimation of the survival function. Predicted ranks were used as risk scores for computing the C-index.
@@
Let $\hat{q}_{i}(t)$ be the predicted risk score of observation with covariates $\mathbf{x}_{i}$ at time $t$. The C-index estimate [HARRELL1996] is given by
$\text{C-index}=\frac{\sum_{i=1}^{N}\sum_{j\neq i}\delta_{i}\>\mathbbm{1}_{\{y_{i}<y_{i}\}}\left(\mathbbm{1}_{\{\hat{q}_{i}(y_{i})>\hat{q}_{j}(y_{i})\}}+\frac{1}{2}\mathbbm{1}_{\{\hat{q}_{i}(y_{i})=\hat{q}_{j}(y_{i})\}}\right)}{\sum_{i=1}^{N}\sum_{j\neq i}\delta_{i}\>\mathbbm{1}_{\{y_{i}<y_{j}\}}}.$
Let $\hat{S}_{i}(t)$ be the predicted survival function of observation with covariates $\mathbf{x}_{i}$ at time $t$. When the predicted risk score is taken to be the negative of the survival function, i.e., $\hat{q}_{i}(t)=-\hat{S}_{i}(t)$, the C-index is referred to as the Antolini’s C-index [Antolini2005] and is found with
$\text{C-index}=\frac{\sum_{i=1}^{N}\sum_{j\neq i}\delta_{i}\>\mathbbm{1}_{\{y_{i}<y_{i}\}}\left(\mathbbm{1}_{\{\hat{S}_{i}(y_{i})<\hat{S}_{j}(y_{i})\}}+\frac{1}{2}\mathbbm{1}_{\{\hat{S}_{i}(y_{i})=\hat{S}_{j}(y_{i})\}}\right)}{\sum_{i=1}^{N}\sum_{j\neq i}\delta_{i}\>\mathbbm{1}_{\{y_{i}<y_{j}\}}}.$
The C-index is obtained using the ConcordanceIndex class from the TorchSurv package [Monod2024].
IPCW Integrated Brier Score.
Let $\hat{S}_{i}(t)$ be the predicted survival function of observation with covariates $\mathbf{x}_{i}$ at time $t$. Let the inverse probability censoring weight (IPCW) at time $t$ be defined as the inverse of the probability of being uncensored, $\xi(t)=1/\hat{C}(t)$, where $\hat{C}(t)$ denotes the Kaplan–Meier estimate of the censoring survival function. Under right censorship, the IPCW Brier score (BS) [Graf1999] at time $t$ is given by
$\text{IPCW BS}(t)=\frac{1}{N}\sum_{i=1}^{N}\xi(y_{i})\mathbbm{1}_{\{y_{i}\leq t,\delta_{i}=1\}}(0-\hat{S}_{i}(t))^{2}+\xi(t)\mathbbm{1}_{\{y_{i}>t\}}(1-\hat{S}_{i}(t))^{2}.$ (A29)
The IBS is the integral of the Brier Score in (A29). The IPCW weights and the IPCW IBS are computed using the get_ipcw function and the BrierScore class from the TorchSurv package [Monod2024].
Distribution Calibration.
D-Calibration [Haider2020] evaluates whether predicted survival probabilities at observed times are uniformly distributed. For an individual $i$, let $\hat{S}_{i}(y_{i})$ be the predicted survival probability at their event or censoring time $y_{i}$. Under perfect calibration, we expect:
$\hat{S}_{i}(y_{i})\sim\text{Uniform}(0,1).$ (A30)
The predicted probabilities are binned into $B$ quantiles, and a histogram is constructed over both event and censored observations. For censored data, the probability is distributed proportionally across bins beyond the censoring time. A chi-squared test compares the resulting histogram to the expected uniform distribution, and the p-value reflects how well the survival model is calibrated. The D-Calibration is obtained using the LifelinesEvaluator.d_calibration() class from the SurvivalEVAL package [qi2024survivaleval].
Kaplan-Meier Calibration.
Kaplan-Meier (KM)-Calibration, as introduced by [Chapfuwa2023], evaluates how well the average predicted survival curve from a model aligns with the empirical KM survival curve. Let $\hat{S}_{\text{avg}}(t)$ denote the model’s average survival probability at time $t$, and $\hat{S}_{\text{KM}}(t)$ the KM estimate. The KM calibration score is defined as the normalized integrated mean squared error (MSE):
$\text{KM-Calibration}=\frac{1}{T_{\max}}\int_{0}^{T_{\max}}\left(\hat{S}_{\text{avg}}(t)-\hat{S}_{\text{KM}}(t)\right)^{2}\,dt.$ (A31)
This score lies in $[0,1]$, where 0 indicates perfect calibration, and values near 0.25 represent uninformative predictions. The KM-Calibration is obtained using the LifelinesEvaluator.km_calibration() class from the SurvivalEVAL package [qi2024survivaleval].
@@
The parameters are initialized so that they match their prior expected values. Specifically, we set $\boldsymbol{\theta}^{(0)}=\mathbf{0}$ and $\phi^{(0)}=\alpha_{0}/\beta_{0}$. The maximization step of the EM algorithm is performed using the L-BFGS-B algorithm. The EM algorithm is considered to have converged when the relative change in the Q-function between consecutive iterations falls below a tolerance threshold of $10^{-6}$ for two successive iterations.